# Localized Schrödinger Bridge Sampler

### v1

@@ 5. Conclusions
The construction of the Schrödinger bridge is fraught with an unfavorable dependency on the dimension $d$. The required number of samples scales for a desired accuracy exponentially on the underlying intrinsic dimensionality of the data [21]. This is exactly the advantage of our proposed localization: Localization requires less samples for a desired accuracy. We have demonstrated numerically the advantage of localization by means of a Gaussian measure for which the inverse covariance matrix has tri-diagonal structure and for a nonlinear close problem arising from the multi-scale Lorenz-96 model. We have also established theoretically that the proposed sampler is stable and geometric ergodic under relatively mild conditions. The stability of our sampler allows for applications to data drawn from a singular measure. This sets it apart from score-generative models which rely on the differentiability of the measure. While this work has focused on overdamped Langevin dynamics as a mean of sampling from a distribution, the methodology generalizes to more general formulations of score-generative and diffusion modeling [7, 17, 19, 22]. Further potential applications include sequential data assimilation [16, 2, 3], feedback particle filter and homotopy methods [23, 15, 14], which are implemented utilizing Schrödinger bridges, and interacting particle sampling methods, which rely on grad-log density estimators [10].

### v3

@@ 6. Conclusions
The construction of the previously proposed Schrödinger bridge sampler [10] is fraught with an unfavorable dependency in the dimension $d$. The required number of samples scales for a desired accuracy exponentially on the underlying intrinsic dimensionality of the data [35]. We have shown here that for data which satisfy conditional independence one can successfully employ localization to express the single Schrödinger bridge problem for $d$-dimensional data to $d$ localized Schrödinger bridge problems of smaller size $d_{\alpha}\ll d$. The localized Schrödinger bridge sampler can be used to generate samples from an unknown prior and readily lends itself to conditional sampling and Bayesian inference.
We have numerically demonstrated the advantage of localization for several examples. We considered a Gaussian distribution for which the inverse covariance matrix has tri-diagonal structure, a bimodal SDE and a conditional sampling problem of determining a closure term in a nonlinear multi-scale system.
We have established theoretically that the proposed sampler is stable and geometric ergodic under relatively mild conditions. The stability of our sampler allows for applications to data drawn from a singular measure which arise when data are concentrated on a lower-dimensional manifold. This sets it apart from score-generative models which rely on Tweedie’s formula and the differentiability of a regularized measure.
We have established several connections with other sampling strategies. The Schrödinger bridge sampler was shown to be closely related to kernel-based denoising. The Schrödinger bridge sampler, however, has the advantage that it can employ a data-aware noising step, which was demonstrated to be advantageous in Section 3.1.1, and can be constructed using a variable bandwidth [10], which is desirable with training data that involve data-sparse regions in the state space. Further, while this work has focused on overdamped Langevin dynamics as a mean of sampling from a distribution, the methodology generalizes to more general formulations of score-generative and diffusion modeling [12, 27, 29, 36] and transformers [30, 26]. We have shown that the conditional mean of a Schrödinger bridge sampler is formally akin to self-attention in transformer architectures and that localization naturally leads to multi-head self attention. It will be interesting to further explore these connections.
The framework of localized Schrödinger bridges lends itself to numerous applications. In particular, we mention here sequential data assimilation [25, 4, 7], feedback particle filter and homotopy methods [37, 24, 23], which are implemented utilizing Schrödinger bridges, and interacting particle sampling methods, which rely on grad-log density estimators such as (1) [17]. Finally, the proposed localized conditional estimator $m_{\rm loc}(x)$ as well as its KDE-based variant (78) could be of independent interest for MMSE denoising [20].